OpenStudy (nottim):

For the following expressions, state the order of the reaction with respect to each especies, the overall order of the reaction, and the units of teh rate constant k: rate= k[HI]^2[O2]/([H+]^(1/2))

OpenStudy (nottim):

hi =2, o2 =1, h+ =1/2. ovERALL= 2.5. iS THAT CORRECT? What about the units?

OpenStudy (nottim):

hi =2, o2 =1, h+ =1/2. ovERALL= 2.5. iS THAT CORRECT? What about the units?

OpenStudy (abhisar):

The rate equation can be written as \(\sf rate= k[HI]^2[O2]/[H+]^{(1/2)}=k[HI]^2[O_2][H^+]^{-0.5}\) So order wrt to [H+]=-0.5 Overall order = 2+1-0.5 = 2.5

OpenStudy (abhisar):

Since, here, K = \(\sf\frac{Rate}{[HI]^2[O_2][H^{+}]^{(-1/2)}}\) Unit of K = \(\frac{Conc.}{Sec} \times \frac{1}{Conc.^{2.5}}=Conc.^{-1.5}/sec=Mol^{-1.5}L^{1.5}/Sec\)

OpenStudy (nottim):

Where did -1.5 come from?

OpenStudy (priyar):

conc^(1-2.5) = con^(-1.5) that's how it came

OpenStudy (priyar):

@NotTim

OpenStudy (nottim):

priyar, i'm sorry, i don't understand.

OpenStudy (nottim):

wait thanks, i figured it out

OpenStudy (priyar):

oh thats great..! u r welcome @NotTim